OpenStudy (bryanna):

1st derivate 2x+(5/2)x^3/2 + 2x^-3 find second and third derivate

OpenStudy (anonymous):

i'm guessing that that's the first derivative. Second derivative = x + (15/4)x^1/2 - 6x^-4 Third derivative = (15/8)x^(-1/2) + 24x^-5

OpenStudy (anonymous):

2x should be 2 (not x) after first derivative

OpenStudy (anonymous):

oh right. made a mistake . thx